you need $11,000 to purchase a used car. Your wealthy uncle is willing to lend you the money as an amortized loan. He would like you to make annual payments for 6 years, with the first payment to be made one year from today. He requires a 9% annual return.

  1. What will be your annual loan payments? Do not round intermediate calculations. Round your answer to the nearest cent.

    $  

  2. How much of your first payment will be applied to interest and to principal repayment? Do not round intermediate calculations. Round your answers to the nearest cent.

    Interest: $  

    Principal repayment: $

3)You borrow $85,000; the annual loan payments are $4,336.64 for 30 years. What interest rate are you being charged? Round your answer to the nearest whole number.
Answer #1

1-a). Annual Loan Payment = [Loan Amount * r] / [1 - (1 + r)-n]

= [$11,000 * 0.09] / [1 - (1 + 0.09)-6]

= $990 / 0.4037 = $2,452.12

1-b). Interest part of the first payment = Outstanding balance at the start of the period * Interest Rate

= $11,000 * 0.09 =$990

Principal Repayment part of the first payment = Annual Loan Payment - Interest part of the first payment

= $2,452.12 - $990 = $1,462.12

3). To find the interest rate charged, we need to put the following values in the financial calculator:

INPUT 30 85,000 -4,336.64 0
TVM N I/Y PV PMT FV
OUTPUT 3

Hence, Interest Rate = 3%